# Multiple commutation relations of the quantum affine algebra $U_{q}(\widehat{\mathfrak{gl}}_{N})$, nested Bethe vector and the Gelfand-Tsetlin basis

### v1

@@ 1 Introduction
Quantum groups [1, 2, 3] are Hopf algebra deformations of universal enveloping algebras of Lie algebras, introduced in connection with integrable systems and the Yang–Baxter equation [4, 5, 6, 7] in the 1980s. Since then, numerous works have been devoted to their structural theory and applications across representation theory, low-dimensional topology, and mathematical physics. Nevertheless, even certain naive questions on quantum groups still remain to be difficult or need conceptual understandings. In this paper, we deal with one of such, the multiple commutation relations. The defining relations of the quantum groups are given as commutation relations between two $L$-operator elements. A naive question is then what are the explicit commutation relations if there are more types and multiple $L$-operator elements. Investigating multiple commutation relations and multiple actions on vectors is a fundamental problem and are important for the study of correlation functions of quantum integrable models [6, 7, 8]. It has also been important for the study of stochastic integrable models [9] and the three-dimensional partition functions [10]. As for higher rank quantum groups and Yangians, significant progress began to emerge recently. See [11, 12, 13] for example. This is also aligned with the progress of algebraic understanding of the nested Bethe vectors and higher rank weight functions [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. The higher rank trigonometric weight functions are extended to the elliptic case in recent years [29, 30, 31, 32].
@@
We denote the symmetric group, the group of all permutations of a set with $n$ elements, by $S_{n}$. We also denote $n$ consecutive occurrences of the number $i$ by $i^{n}$. This type of notation appears in the description of the (dual) orthonormal basis.
@@
$J_{1}\cup\cdots\cup J_{N}=\{1,\dots,n\},\ \ \ J_{j}\cap J_{k}=\phi\ (j\neq k).$ (1.3)
Typical (sub)sets that appear are $\overline{w}=\{w_{1},\dots,w_{n}\}$ and $\overline{w}_{J_{j}}=\{w_{k}\ |\ k\in J_{j}\}$ ($j=1,\dots,N$).
@@ 2 Quantum affine algebra $U_{q}(\widehat{\mathfrak{gl}}_{N})$ and basic partition functions
The non-zero $R$-matrix elements $R(u,v)]_{ij}^{k\ell}$ satisfy $k=i$ and $j=\ell$ if $i\neq\ell$ for example, which are more restrictive than the so-called ice rule: $R(u,v)]_{ij}^{k\ell}$=0 unless $i+j=k+\ell$.
@@
Here, $T_{1}(u)=T(u)\otimes 1$, $T_{2}(v)=1\otimes T(v)$ (and similarly for $\bar{T}$). In this paper, we refer to $T(u)$ as the L-operator and $T_{ij}(u)$, $i,j=1,\dots,N$ as the L-operator elements (also called the monodromy matrix and the monodromy matrix elements in the literature).
@@
where $|\overline{u}|=|\overline{v}|$. The domain wall boundary partition functions are symmetric with respect to the set of variables $\overline{u}$ as well as $\overline{v}$, due to the Yang-Baxter relation.
@@ 3 Multiple commutation relations
$e_{1^{|\overline{u}^{1}|},2^{|\overline{u}^{2}|},\dots,(N-1)^{|\overline{u}^{N-1}|}}^{*}T_{NN}(\overline{v}^{N};\overline{v}^{1}_{0},\overline{v}^{2}_{0},\dots,\overline{v}^{N-1}_{0})$. Recall that for $\ell\neq N$, $R_{Nj}^{k\ell}=0$ unless $j=\ell$, $k=N$. Using this property, we can see the action is uniquely determined and only the matrix elements $R_{Nj}^{Nj}$, $j=1,\dots,N-1$ appear, and we get
@@
Next, consider $e_{1^{|\overline{u}^{1}|},2^{|\overline{u}^{2}|},\dots,(N-1)^{|\overline{u}^{N-1}|}}^{*}T_{NN-1}(\overline{v}^{N-1};\overline{v}^{1}_{0},\overline{v}^{2}_{0},\dots,\overline{v}^{N-1}_{0})$. By a similar observation, we note the action of $T_{NN-1}(\overline{v}^{N-1};\overline{v}^{1}_{0},\overline{v}^{2}_{0},\dots,\overline{v}^{N-1}_{0})$ on the dual quantum spaces except the last $|\overline{u}|^{N-1}$ spaces are uniquely determined, and the matrix elements $R_{N-1,j}^{N-1,j}$, $j=1,\dots,N-2$ appear which give the factor $\prod_{k=1}^{N-2}(\overline{v}^{N-1}-\overline{v}^{k}_{0})$. As for the last $|\overline{u}^{N-1}|$ quantum spaces, we need to consider $e_{(N-1)^{|\overline{u}^{N-1}|}}^{*}T_{N,N-1}(\overline{v}^{N-1};\overline{v}^{N-1}_{0})$. By a simple observation, we conclude the resulting state must be $e_{N^{|\overline{u}^{N-1}|}}^{*}$ (the resulting state must be a linear combination of $e_{i_{1},i_{2},\dots,i_{|\overline{u}^{1}|+\cdots+|\overline{u}^{N-1}|}}^{*}$ where $i_{j}=N-1,N$ since $T_{N-1,N}$ is constructed from the trigonometric $R$-matrix, and the conservation of colors lead to $i_{j}=N-1$ for all $j$), and we get
@@
from which one can see only summands satisfying $\overline{v}^{N}\cap\overline{v}^{1}_{0}=\overline{v}^{N}\cap\overline{v}^{2}_{0}=\cdots=\overline{v}^{N}\cap\overline{v}^{N-1}_{0}=\phi$ survive, i.e., $\overline{v}^{N}$ must be $\overline{v}^{N}=\overline{v}^{N}_{0}$. Next, one can see nonzero summands should satisfy $\overline{v}^{N-1}\cap\overline{v}^{1}_{0}=\cdots=\overline{v}^{N-1}\cap\overline{v}^{N-2}_{0}=\phi$, and combining with $\overline{v}^{N}=\overline{v}^{N}_{0}$ one concludes $\overline{v}^{N-1}=\overline{v}^{N-1}_{0}$. Repeating this argument, we find only the summand corresponding to $\overline{v}^{j}=\overline{v}^{j}_{0}$, $j=1,\dots,N$ survives. We can also show
@@
Figure 9: Graphical description of (3.14). In the left panel, we observe that in lower left-most intersection point the spaces share the same spectral parameter $v_{1}$, which reduces the possible $R$-matrix elements to just a single choice. This then propagates throughout the first row and column. The same is repeated for the second row and column, and so on, giving just a single term for the whole expression (right panel).
@@
Figure 11: A graphical description of the enlarged partition function $K$. In the rightmost columns, the only possible color is $N$, which proves Lemma 3.4.
@@
We next note that the following relation holds. This is essentially the same as a Proposition in Borodin-Wheeler [24, Proposition 7.1.1]. We use the standard trigonometric $R$-matrix.
@@
This follows as a specific case $\bm{I}=(1^{|\overline{u}^{1}|},2^{|\overline{u}^{2}|},\dots,(N-1)^{|\overline{u}^{N-1}|},N^{|\overline{u}^{N}|})$ of a more generic relation between partition functions and Theorem 2.2. For the description, we introduce a partition $J=(J_{1},\dots,J_{N})$ such that
$J_{1}\cup\cdots\cup J_{N}=\{1,\dots,n\},\ \ \ J_{j}\cap J_{k}=\phi\ (j\neq k).$ (3.24)
Let $\overline{w}=\{w_{1},\dots,w_{n}\}$ and $\overline{w}_{J_{j}}=\{w_{k}\ |\ k\in J_{j}\}$ ($j=1,\dots,N$). Note $\overline{w}=\overline{w}_{J_{1}}\cup\cdots\cup\overline{w}_{J_{N}}$.
@@
The trick we use is essentially the same with the one given in [24, Proposition 7.1.1]. Observe that the ordering of the spaces in each layer is arbitrary – the input and output coordinate is the same in each layer, so reordering can be accomplished by insertion of an $R$-matrix and application of the Yang-Baxter equation. We thus choose the lexicographic ordering according to the parameter label from left to right and top to bottom in each case.
@@
Finally, as the remaining coordinates on the right-hand side are all equal to $j+1$, this part of the partition function is also fully determined, with factor $(q\overline{w}_{J_{1}\cup\cdots\cup{J_{j}}}-q^{-1}\,\overline{w}_{J_{j+1}})$. Combining this with (3.30), we have an overall factor of
@@
As a final remark of this section, let us present another form for the case $N=2$, recovering another presentation using the Izergin-Korepin determinants [7, (7.62), (7.63)]. The case $N=2$ of Proposition 3.2 is
@@ 4 A construction the Gelfand-Tsetlin basis for the vector representation
The relation (3.25) implies that $T_{N1}(\overline{w}_{J_{1}};\overline{w})T_{N2}(\overline{w}_{J_{2}};\overline{w})\ \cdots T_{NN-1}(\overline{w}_{J_{N-1}};\overline{w})e_{N^{n}}$ for all $J$ give rise to a construction of the Gelfand-Tsetlin basis for the case of the tensor product of vector representations, a basis which simultaneously diagonalizes the Gelfand-Tsetlin subalgebra. There are two known main constructions of the Gelfand-Tsetlin basis, one due to Nazarov-Tarasov [27] and another due to Molev [28]. We briefly review Molev’s construction in the Appendix, which shows that $T_{21}(w_{J_{1}};\overline{w})T_{32}(w_{J_{1}}\cup w_{J_{2}};\overline{w})\cdots T_{N,N-1}(w_{J_{1}}\cup\cdots\cup w_{J_{N-1}};\overline{w})e_{N^{n}}$ for all $J$ give rise to a construction of the Gelfand-Tsetlin basis. However, it seems not easy to show directly that $T_{N1}(\overline{w}_{J_{1}};\overline{w})T_{N2}(\overline{w}_{J_{2}};\overline{w})\cdots T_{NN-1}(\overline{w}_{J_{N-1}};\overline{w})e_{N^{n}}$ for $N>2$ diagonalizes the Gelfand-Tsetlin subalgebra. We take an indirect approach and derive the following relations between $T_{N1}(\overline{w}_{J_{1}};\overline{w})T_{N2}(\overline{w}_{J_{2}};\overline{w})\cdots T_{NN-1}(\overline{w}_{J_{N-1}};\overline{w})e_{N^{n}}$ and $T_{21}(w_{J_{1}};\overline{w})T_{32}(w_{J_{1}}\cup w_{J_{2}};\overline{w})\cdots T_{N,N-1}(w_{J_{1}}\cup\cdots\cup w_{J_{N-1}};\overline{w})e_{N^{n}}$.
@@
Two expressions in [11, Proposition 3.1] for the universal Bethe vector lead to the following identity.
@@
[11, Proposition 3.1] The following relation holds:
@@
$\Psi$ $=(\overline{t}^{1}-\overline{w})\prod_{\ell=2}^{N-1}\left(\overline{t}^{\ell}-\overline{t}^{\ell-1}\right)\widehat{B}(\bar{t}^{\,1},\bar{t}^{\,2},\dots,\bar{t}^{\,N-1})e_{1^{n}}.$ (4.12)
@@
In the next two subsections, we provide the details of specializing $\Psi$ and $\widetilde{\Psi}$ to get (4.19) and (4.23).

### v2

@@ 1 Introduction
Quantum groups [1, 2, 3] are Hopf algebra deformations of universal enveloping algebras of Lie algebras, introduced in connection with integrable systems and the Yang–Baxter equation [4, 5, 6, 7, 8] in the 1980s. Since then, numerous works have been devoted to their structural theory and applications across representation theory, low-dimensional topology, and mathematical physics. Nevertheless, even certain naive questions on quantum groups still remain to be difficult or need conceptual understandings. In this paper, we deal with one of such, the multiple commutation relations. The defining relations of the quantum groups are given as commutation relations between two $L$-operator elements. A naive question is then what are the explicit commutation relations if there are more types and multiple $L$-operator elements. Investigating multiple commutation relations and multiple actions on vectors is a fundamental problem and are important for the study of correlation functions of quantum integrable models [7, 8, 9]. It has also been important for the study of stochastic integrable models [10] and the three-dimensional partition functions [11]. As for higher rank quantum groups and Yangians, significant progress began to emerge recently. See [12, 13, 14] for example. This is also aligned with the progress of algebraic understanding of the nested Bethe vectors and higher rank weight functions [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. The higher rank trigonometric weight functions are extended to the elliptic case in recent years [31, 32, 33, 34].
@@
We denote the symmetric group, the group of all permutations of a set with $n$ elements, by $S_{n}$. We also denote $m$ consecutive occurrences of the number $i$ by $i^{m}$. This type of notation appears in the description of the (dual) orthonormal basis, for example $e_{i^{m}}=\underbrace{e_{i}\otimes\cdots\otimes e_{i}}_{m\text{ times}}$.
@@
$J_{1}\cup\cdots\cup J_{N}=\{1,\dots,n\},\ \ \ J_{j}\cap J_{k}=\varnothing\ (j\neq k).$ (1.3)
The $J_{j}$ are permitted to be empty. Typical (sub)sets that appear are $\overline{w}=\{w_{1},\dots,w_{n}\}$ and $\overline{w}_{J_{j}}=\{w_{k}\ |\ k\in J_{j}\}$ ($j=1,\dots,N$).
@@ 2 Quantum affine algebra $U_{q}(\widehat{\mathfrak{gl}}_{N})$ and basic partition functions
The $R$-matrix elements $R(u,v)]_{ij}^{k\ell}$ satisfy
$R(u,v)]_{ij}^{k\ell}=0\qquad\text{unless }(i=k,\,j=\ell)\text{ or }(i=\ell,\,j=k).$ (2.6)
This is more restrictive than the ice rule, where $R(u,v)]_{ij}^{k\ell}$=0 unless $i+j=k+\ell$, which is obeyed by the $R$-matrices of higher spin representations of $U_{q}(\widehat{\mathfrak{sl}}_{2})$. We will refer to the property (2.6) as color conservation.
@@
Here, $T_{1}(u)=T(u)\otimes 1$, $T_{2}(v)=1\otimes T(v)$ (and similarly for $\bar{T}$). The above definition is for the case where the central charge is equal to zero; for the generic central charge case see [39, 40] for example. In this paper, we refer to $T(u)$ as the L-operator and $T_{ij}(u)$, $i,j=1,\dots,N$ as the L-operator elements (also called the monodromy matrix and the monodromy matrix elements in the literature).
@@
where $|\overline{u}|=|\overline{v}|$. Here we recall $e_{2^{|\overline{u}|}}$ refers to the vector $\underbrace{e_{2}\otimes\dots\otimes e_{2}}_{|\overline{u}|\text{ times}}$. The domain wall boundary partition functions are symmetric with respect to the set of variables $\overline{u}$ as well as $\overline{v}$, due to the Yang-Baxter relation.
@@ 3 Multiple commutation relations
$e_{1^{|\overline{u}^{1}|},2^{|\overline{u}^{2}|},\dots,(N-1)^{|\overline{u}^{N-1}|}}^{*}T_{NN}(\overline{v}^{N};\overline{v}^{1}_{0},\overline{v}^{2}_{0},\dots,\overline{v}^{N-1}_{0})$. Applying (2.6), for $\ell\neq N$, $R_{Nj}^{k\ell}=0$ unless $j=\ell$, $k=N$. Using this property, we can see the action is uniquely determined and only the matrix elements $R_{Nj}^{Nj}$, $j=1,\dots,N-1$ appear, and we get
@@
Next, consider $e_{1^{|\overline{u}^{1}|},2^{|\overline{u}^{2}|},\dots,(N-1)^{|\overline{u}^{N-1}|}}^{*}T_{NN-1}(\overline{v}^{N-1};\overline{v}^{1}_{0},\overline{v}^{2}_{0},\dots,\overline{v}^{N-1}_{0})$. By a similar observation, we note the action of $T_{NN-1}(\overline{v}^{N-1};\overline{v}^{1}_{0},\overline{v}^{2}_{0},\dots,\overline{v}^{N-1}_{0})$ on the dual quantum spaces except the last $|\overline{u}|^{N-1}$ spaces are uniquely determined, and the matrix elements $R_{N-1,j}^{N-1,j}$, $j=1,\dots,N-2$ appear which give the factor $\prod_{k=1}^{N-2}(\overline{v}^{N-1}-\overline{v}^{k}_{0})$. As for the last $|\overline{u}^{N-1}|$ quantum spaces, we need to consider $e_{(N-1)^{|\overline{u}^{N-1}|}}^{*}T_{N,N-1}(\overline{v}^{N-1};\overline{v}^{N-1}_{0})$. By a simple observation, we conclude the resulting state must be $e_{N^{|\overline{u}^{N-1}|}}^{*}$ (the resulting state must be a linear combination of $e_{i_{1},i_{2},\dots,i_{|\overline{u}^{1}|+\cdots+|\overline{u}^{N-1}|}}^{*}$ where $i_{j}=N-1,N$ since $T_{N-1,N}$ is constructed from the trigonometric $R$-matrix, and the conservation of colors (2.6) lead to $i_{j}=N-1$ for all $j$), and we get
@@
from which one can see only summands satisfying $\overline{v}^{N}\cap\overline{v}^{1}_{0}=\overline{v}^{N}\cap\overline{v}^{2}_{0}=\cdots=\overline{v}^{N}\cap\overline{v}^{N-1}_{0}=\varnothing$ survive, i.e., $\overline{v}^{N}$ must be $\overline{v}^{N}=\overline{v}^{N}_{0}$. Next, one can see nonzero summands should satisfy $\overline{v}^{N-1}\cap\overline{v}^{1}_{0}=\cdots=\overline{v}^{N-1}\cap\overline{v}^{N-2}_{0}=\varnothing$, and combining with $\overline{v}^{N}=\overline{v}^{N}_{0}$ one concludes $\overline{v}^{N-1}=\overline{v}^{N-1}_{0}$. Repeating this argument, we find only the summand corresponding to $\overline{v}^{j}=\overline{v}^{j}_{0}$, $j=1,\dots,N$ survives. We can also show
@@
Figure 9: Graphical description of (3.14). In the left panel, we observe that in lower left-most intersection point the spaces share the same spectral parameter $v_{1}$, which reduces the possible $R$-matrix elements to just a single choice. This then propagates throughout the first row and column due to color conservation (2.6). The same is repeated for the second row and column, and so on, giving just a single term for the whole expression (right panel).
@@
Figure 11: A graphical description of the enlarged partition function $K$. In the rightmost columns, the only possible color is $N$ due to color conservation (2.6), which proves Lemma 3.4.
@@
We next note that the following relation holds. This is essentially the same as a Proposition in Borodin-Wheeler [26, Proposition 7.1.1]. We use the standard trigonometric $R$-matrix.
@@
This follows as a specific case $\bm{I}=(1^{|\overline{u}^{1}|},2^{|\overline{u}^{2}|},\dots,(N-1)^{|\overline{u}^{N-1}|},N^{|\overline{u}^{N}|})$ of a more generic relation between partition functions and Theorem 2.2. For the description, we introduce a partition $J=(J_{1},\dots,J_{N})$ as in (1.3). Let $\overline{w}=\{w_{1},\dots,w_{n}\}$ and $\overline{w}_{J_{j}}=\{w_{k}\ |\ k\in J_{j}\}$ ($j=1,\dots,N$). Note $\overline{w}=\overline{w}_{J_{1}}\cup\cdots\cup\overline{w}_{J_{N}}$.
@@
The trick we use is essentially the same with the one given in [26, Proposition 7.1.1]. Observe that the ordering of the spaces in each layer is arbitrary – the input and output coordinate is the same in each layer, so reordering can be accomplished by insertion of an $R$-matrix and application of the Yang-Baxter equation. We thus choose the lexicographic ordering according to the parameter label from left to right and top to bottom in each case.
@@
Finally, as the remaining coordinates on the right-hand side are all equal to $j+1$, this part of the partition function is also fully determined by the conservation of colors (2.6). Each $R$-matrix element that appears is equal to $R(u,v)]_{j+1,j+1}^{j+1,j+1}$ for some $u,v$, resulting in factor $(q\overline{w}_{J_{1}\cup\cdots\cup{J_{j}}}-q^{-1}\,\overline{w}_{J_{j+1}})$. Combining this with (3.29), we have an overall factor of
@@
As a final remark of this section, let us present another form for the case $N=2$, recovering another presentation using the Izergin-Korepin determinants [8, (7.62), (7.63)]. The case $N=2$ of Proposition 3.2 is
@@ 4 A construction the Gelfand-Tsetlin basis for the vector representation
The relation (3.24) implies that $T_{N1}(\overline{w}_{J_{1}};\overline{w})T_{N2}(\overline{w}_{J_{2}};\overline{w})\ \cdots T_{NN-1}(\overline{w}_{J_{N-1}};\overline{w})e_{N^{n}}$ for all $J$ give rise to a construction of the Gelfand-Tsetlin basis for the case of the tensor product of vector representations, a basis which simultaneously diagonalizes the Gelfand-Tsetlin subalgebra. There are two known main constructions of the Gelfand-Tsetlin basis, one due to Nazarov-Tarasov [29] and another due to Molev [30]. We briefly review Molev’s construction in the Appendix, which shows that $T_{21}(w_{J_{1}};\overline{w})T_{32}(w_{J_{1}}\cup w_{J_{2}};\overline{w})\cdots T_{N,N-1}(w_{J_{1}}\cup\cdots\cup w_{J_{N-1}};\overline{w})e_{N^{n}}$ for all $J$ give rise to a construction of the Gelfand-Tsetlin basis. However, it seems not easy to show directly that $T_{N1}(\overline{w}_{J_{1}};\overline{w})T_{N2}(\overline{w}_{J_{2}};\overline{w})\cdots T_{NN-1}(\overline{w}_{J_{N-1}};\overline{w})e_{N^{n}}$ for $N>2$ diagonalizes the Gelfand-Tsetlin subalgebra. We take an indirect approach and derive the following relations between
$T_{N1}(\overline{w}_{J_{1}};\overline{w})T_{N2}(\overline{w}_{J_{2}};\overline{w})$ $\cdots T_{NN-1}(\overline{w}_{J_{N-1}};\overline{w})e_{N^{n}}$ and $T_{21}(w_{J_{1}};\overline{w})T_{32}(w_{J_{1}}\cup w_{J_{2}};\overline{w})\cdots T_{N,N-1}(w_{J_{1}}\cup\cdots\cup w_{J_{N-1}};\overline{w})e_{N^{n}}$.
@@
Two expressions in [12, Proposition 3.1] for the universal Bethe vector lead to the following identity.
@@
[12, Proposition 3.1] The following relation holds:
@@
$\widetilde{\Psi}$ $=(\overline{t}^{1}-\overline{w})\prod_{\ell=2}^{N-1}\left(\overline{t}^{\ell}-\overline{t}^{\ell-1}\right)\widehat{B}(\bar{t}^{\,1},\bar{t}^{\,2},\dots,\bar{t}^{\,N-1})e_{1^{n}}.$ (4.12)
@@
In subsections 4.2 and 4.3, we provide the details of specializing $\Psi$ and $\widetilde{\Psi}$ to get (4.20) and (4.24).
4.1 Relation to the trace formula
For the $U_{q}(\widehat{{\mathfrak{gl}}}_{3})$ case, the Tarasov-Varchenko trace formula for the Bethe vector [22] is given by
[…]
Figure 18: A pictorial representation of the trace formula for $U_{q}(\widehat{{\mathfrak{gl}}}_{3})$.
[…]
Figure 19: The lower-right grid becomes fully determined using (2.6).
As a result, the trace formula is equivalent to Figure 20.
Figure 20: The result of applying (2.6) to Figure 19
Specifying the top boundary of this figure essentially means taking a component of the trace formula of the nested Bethe vector and corresponds to $N=3$ case of Figure 5 in the previous section. As in the previous section, further specializing $\{v_{1},\dots v_{b}\}\to\overline{w}_{J_{1}}$ and $\{u_{1},\dots u_{a}\}\to\overline{w}_{J_{1}}\cup\overline{w}_{J_{2}}$ gives
$T_{31}(\overline{w}_{J_{1}};\overline{w})T_{32}(\overline{w}_{J_{2}};\overline{w})e_{3^{p}},$
which agrees with the right-hand side of (4.1) in the case $N=3$ up to an overall factor.